OpenStudy (anonymous):

Find the values of x for which the series converges: ((summation)(n=1 to infinity)) (x-1)^n/3^n

OpenStudy (amistre64):

lim of an/a(n-1) as n goes inf

OpenStudy (amistre64):

\[\lim \frac{(x-1)^n}{3^n}\frac{3^{n-1}}{(x-1)^{n-1}}\] \[\lim \frac{(x-1)}{3}\] \[|\frac{x-1}{3}|\lim 1\]

OpenStudy (anonymous):

I thought it was n+1??

OpenStudy (amistre64):

\[|\frac{x-1}{3}|<1\] \[|x-1|<3\] -2 < x < 4

OpenStudy (amistre64):

its simply next/now how you wanna represent that is up to you

OpenStudy (anonymous):

Thanks!! I understand

OpenStudy (amistre64):

yw :)